OpenStudy (anonymous):

B is the midpoint of A and C. If BC=3x2^2, and AC=54, find the value of x. Show all work to receive credit.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3x^2 or 3x*2^2

OpenStudy (anonymous):

I dont quite get this, explain more please.

OpenStudy (anonymous):

No, i mean the first part , what is BC: 3x^2 or 3x*2^2

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

or \[3x ^{2}\]

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

But if I substitute 3 for x it wont equal 54?

OpenStudy (anonymous):

bc is mid of ac so bc=ac/2= 27 not 54

OpenStudy (anonymous):

ac= 54 and bc is mid of ac so bc - 54/2-27 and acc to eq bc=3x^2=27 divide bothsides by 3 (3x^2)/3=27/3 x^2=9 applying square root x=3

OpenStudy (anonymous):

Thank you.